OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 8/x and g(x) = 8/x

OpenStudy (neer2890):

f(x)=8/x g(x)=8/x. f(g(x))=f(8/x) put 8/x inplace of x in f(x) \[f(8/x)=\frac{ 8 }{ \frac{ 8 }{ x } }=8*\frac{ x }{ 8 }=x\] similarly find g(f(x))=g(8/x)

OpenStudy (anonymous):

I understand now. Thank you!

OpenStudy (neer2890):

you're welcome...:)